OpenStudy (anonymous):

find the solution to the system of equations by substitution x=6 x+y+4=0

OpenStudy (anonymous):

\[x+y+4=0 6+y+4=0 y=- 10\]

OpenStudy (babynini):

@funds Well, you know x = 6, so just plug that into your second equation and solve! Your equation: 6+y+4=0 Do you know how to do it from there? :)